OpenStudy (anonymous):

one solution , no solution , many solution ar infinitely many solution 4x - 7y = 15-8x + 14y = -30

Directrix (directrix):

Is this --> 4x - 7y = 15-8x + 14y = -30 supposed to be "Solve the system of equations" 4x - 7y = 15 -8x + 14y = -30

OpenStudy (anonymous):

yes

Directrix (directrix):

Take the first equation --> 4x - 7y = 15 and multiply each term by 2. Post what you get.

OpenStudy (anonymous):

i get the same as the second equation so it a infinitely many solution

Directrix (directrix):

So, the two given equations are the same equation. I agree with you on the infinitely many solutions - a line has infinitely many points.

OpenStudy (anonymous):

thank u very much @Directrix

Directrix (directrix):

You did the work. Glad I was along for the solution.

OpenStudy (anonymous):

lol